OpenStudy (anonymous):

Inverse of this function \[y=x^3-6x^2+12x \] one medal will be given to correct answer

OpenStudy (lgbasallote):

by "inverse" do you mean something like f^-1 (x)?

OpenStudy (anonymous):

yeah...

OpenStudy (rsadhvika):

(x-8)^1/3 + 2

OpenStudy (anonymous):

can by written as \[\Large y=(x-2)^3+8\]

OpenStudy (anonymous):

now \[\Large y-8=(x-2)^3\]

OpenStudy (anonymous):

replace each x with y and each y with x

OpenStudy (anonymous):

\[\Large x-8=(y-2)^3\]

OpenStudy (anonymous):

and replace it with f^-1(x)

OpenStudy (anonymous):

sorry typo!

OpenStudy (anonymous):

\[\Large (x-2)^{1/3}+2=f^{-1}(x)\]

OpenStudy (anonymous):

\[f^{-1}(x)=\sqrt[3]{x-8}+2 \]

OpenStudy (anonymous):

u made a little mistake ..........:D

OpenStudy (anonymous):

yes i did it again !

OpenStudy (anonymous):

m quite notorious for typo mistakes :P

OpenStudy (anonymous):

lol.....:)

OpenStudy (anonymous):

\[\Large \sqrt[3]{x-8}+2=f^{-1}(x)\] just wanted to correct it ! :)